OpenStudy (isabel):

What is the product in lowest terms? -8/9x5/6

alones (alones):

\[\frac{ -8 }{ 9 }*\frac{ 5 }{ 6}\] \[\frac{ \frac{ -40 }{ 9} }{ 6 }\]

alones (alones):

Divide that and you'll geet your final answer .-.